# Hypothesis Generation with Large Language Models

### v1

@@ 4.2 Generalization of the Generated Hypotheses
We find that the hypotheses generated by HypoGeniC generalize across models (Table 2). Generally, we find Claude-2.1 and Mixtral to be better at inference. Thus, substituting the inference model with them lead to better performance for hypothesis generated with GPT-3.5-turbo. Subsituting Claude-2.1 and Mixtral as each other’s inference model lead to small changes in performance. On Shoe sales, the performance remains high (>90%) regardless of the inference model used.
@@
Table 3: Performance on out-of-distribution deceptive reviews.
@@
These results suggest that the hypotheses generated by HypoGeniC are generalizable across different LLMs, which somewhat contracts the claim in Qiu et al. (2024) that LLMs cannot reliably interpret the hypotheses. We suspect that the reason may be that our tasks only rely on natural language, while their tasks rely on notions of worlds and can be fed into symbolic interpreters.
@@
Table 3 presents an overview for the OOD deceptive review dataset. This dataset differs from Deceptive reviews by including reviews from four cities sourced from different websites (Li et al., 2013). We find that HypoGeniC outperforms few-shot learning by an average of 19.1%. Despite the distribution shift, HypoGeniC surprisingly increases accuracy from Deceptive reviews by an average of 3.3%, suggesting our hypotheses generalize well to this OOD dataset. Claude-2.1 remains the best performing model. In comparison, the performance of RoBERTa drops by 11%. As a result, HypoGeniC with Claude-2.1 outperforms RoBERTa by 1.7%, demonstrating the robustness of hypothesis-based inference. Refer to § C.3 for more details.
@@ 4.3 Qualitative Analysis
For the synthetic dataset, all models are able to find the true underlying hypothesis for Shoe sales: “customers tend to buy shoes that match the color of their shirt.” For the real-world datasets, we compare our hypotheses with findings from the literature. We confirm the validity of some of our hypotheses and discover new insights about the tasks that previous studies did not touch upon. We show a few examples in Table 4, and the full list of hypotheses can be found in Appendix D.
Our hypotheses confirm useful features in existing literature. For Deceptive reviews, we find that deceptive reviews are more likely to be emotional, use superlatives, or contain information that could not have been directly experienced. Similar findings are also found by previous studies on Deceptive reviews (Lai et al., 2020; Anderson & Simester, 2014; Ott et al., 2011; Li et al., 2014). For Tweet popularity, we discover that tweets that are short and concise, with specific or relevant hashtags, or with emotional tones are more likely to be retweeted more, aligning with prior studies (Tan et al., 2014; Gligorić et al., 2019). For we find that revealing something new or using vivid language and imagery can drive engagement from readers to click on headlines. Previous studies also find these rules apply to online news headlines (Banerjee & Urminsky, 2021; Sadoski et al., 2000).
We also discover new insights with our generated hypotheses. For the Deceptive reviews dataset, truthful reviews could mention the reviewer’s purpose for staying at the hotel (e.g., business trip, vacation), but deceptive ones tend not to have this information. For we find that headlines that frame the content in a personal or relatable way are clicked more. For Tweet popularity, tweets that mention influential individuals or organizations are more likely to be retweeted.
Intriguingly, one of our hypotheses contradicts a feature engineering result. Ott et al. (2011) find that the token “future” is associated with deceptive reviews, while one of our hypotheses says that mentions of “past experiences or future travel plans” are indicative of truthfulness. This discrepancy is interesting, because the context for the token “future” is unclear. It could be mentioned in the context of future plans but could also be mentioned as a complaint about “never going to stay at the hotel in the future.” Feature engineering is limited due to the contextual ambiguity, whereas our generated hypotheses and their interpretation by LLMs overcome such limitations.
Our automatic evaluation of hypothesis quality also reflects negative findings. Given mixed evidence from previous literature on the effect of “reading ease” on headline clicks, Banerjee & Urminsky (2021) finds that reading ease negatively impacts click-through rates in hrough careful feature engineering. Consistent with this result, we found that the hypotheses that claim “straightforward” and “clear” writing to be indicative of higher click-through rates have relatively lower accuracies during training.
@@
Deceptive reviews
Deceptive reviews contain more emotional terms.
Li et al. (2014)
Truthful reviews would mention weddings or special occasions.
Using vivid language and imagery helps.
Banerjee & Urminsky (2021)
Headlines that frame the content in a personal or relatable way are clicked more.
Tweet popularity
Tweets with emotional tones are retweeted more.
Tan et al. (2014)
Mentioning influential individuals or organizations leads to more retweets.
@@ 5 Additional Related Work
Concept/pattern discovery. In addition to Qiu et al. (2024) and Zhong et al. (2023) discussed in the introduction, other studies have worked along similar lines (Wang et al., 2023b; Singh et al., 2023; Piriyakulkij & Ellis, 2024). For example, similar to Qiu et al. (2024), Tenenbaum et al. (2011) is motivated by human inductive reasoning and examines concept induction in synthetic settings. Ellis et al. (2020) further learns to program concepts. Similar to Zhong et al. (2023), Pham et al. (2024) generate and refine a list of topics to achieve interpretable topic modeling for open-ended exploration. Relatedly, Honovich et al. (2022) explore the deduction of task description from examples. Our work, in contrast, focuses on hypothesis generation between the input and the label for real-world challenging tasks and uses a UCB-style reward to propose novel algorithms.
@@
Reasoning with LLMs. Although it is not our primary goal, our results show that hypothesis-based classifiers can outperform few-shot prompting. As hypotheses may be viewed as a form of reasoning, it is related to reasoning with LLMs (Wei et al., 2022; Wang et al., 2023a, i.a.). In particular, our work differs from chain-of-thought reasoning because no predefined reasoning structure is available. Moreover, an important distinction between reasoning and hypothesis generation is that the former leverages established reasoning, while the latter requires both proposition and verification of the hypotheses, with the goal of discovering unknown knowledge.
@@
LLMs for (social) sciences. Increasing attention has been brought to the use of LLMs in social science research (Ziems et al., 2024; Kim & Lee, 2023, i.a.). Our experiments demonstrate the potential of LLMs in generating hypotheses for social science research to discover unknown knowledge in the data. Furthermore, our approach can be extended to natural sciences for general scientific discovery.
6 Conclusion
@@
In this work, we propose HypoGeniC, a novel method that leverages LLMs to generate hypotheses with the goal of discovering unknown knowledge. With HypoGeniC, we are not only are able to generate human-interpretable hypotheses but also achieve better predictive performance against competitive baselines and even oracles. Furthermore, our method can generalize well with different models and datasets, including open models. Notably, with our generated hypotheses, we uncover new insights in real-world tasks that are widely studied in social sciences. HypoGeniC can be directly applied to natural language related tasks in social sciences. We encourage future work to explore hypothesis generation that requires additional modalities and/or leverages existing literature.
@@ 7 Limitations
In this paper, we aim to provide a robust framework for hypothesis generation, as opposed to focusing on the optimization of results. Thus, we did not perform an extensive hyperparameter search with the generation portion of HypoGeniC. We did not adjust the value of $k$, which determines $\mathcal{H}_{\mathrm{top}}$ in Algorithm 1 to maintain efficiency. Additionatlly, we only considered the effect of using a hypothesis bank size of $3$ and $20$ to only test using an extremely small hypothesis bank size and a large one. The ideal hypothesis bank size may require further investigation. Finally, we only tested the size of our wrong example bank $w_{max}$ as $10$ to strike a balance between context window sizes and generation of good quality hypotheses. We believe that a more thorough hyperparameter search could improve the performance of our methodology.
Additionally, HypoGeniC has high latency, specifically when using inference methods that require multiple prompts. For example, the filter and weighted vote inference policy requires iterating through the top hypotheses to determine relevance and then performing inference if it is relevant. For single-step adaptive inference and best accuracy hypothesis, however, HypoGeniC is efficient. We also note that the main bottleneck with performing inference lies with performing API calls to the LLM, which is a limitation not directly related to HypoGeniC. Given that we request reasoning for all inference prompts, the procedure can be time-consuming.
@@ A.4 Retweeted More
Appendix B Implementation Details

### v2

@@ 4.2 Generalization of the Generated Hypotheses
We find that the hypotheses generated by HypoGeniC generalize across models (Table 2).
Generally, we find Claude-2.1 and Mixtral to be better at inference. Thus, substituting the inference model with them lead to better performance for hypothesis generated with GPT-3.5-turbo. Substituting Claude-2.1 and Mixtral as each other’s inference model lead to small changes in performance. On Shoe sales, the performance remains high for any inference model used.
@@
Table 3: Performance on OOD deceptive reviews.
@@
These results suggest that the hypotheses generated by HypoGeniC are generalizable across different LLMs, which somewhat contradicts the claim in Qiu et al. (2024) that LLMs cannot reliably interpret the hypotheses. We suspect that the reason is that our tasks only rely on natural language, while their tasks rely on notions of worlds and are fed into symbolic interpreters.
@@
Table 3 presents an overview for the OOD deceptive review dataset. This dataset differs from Deceptive reviews by including reviews from four cities sourced from different websites (Li et al., 2013). We find that HypoGeniC outperforms few-shot learning by an average of 19.1%. Despite the distribution shift, HypoGeniC surprisingly increases accuracy from Deceptive reviews by an average of 3.3%, suggesting our hypotheses generalize well to this OOD dataset. Claude-2.1 remains the best performing model. In comparison, the performance of RoBERTa drops by 11%, and Llama-2-7B drops by 10%. As a result, HypoGeniC with Claude-2.1 outperforms RoBERTa by 1.7%, demonstrating the robustness of hypothesis-based inference. Refer to § C.3 for more details.
@@ 4.3 Qualitative Analysis
For the synthetic dataset, all models are able to find the true underlying hypothesis for Shoe sales: “customers tend to buy shoes that match the color of their shirt.” For the real-world datasets, we search for studies on these datasets on Google Scholar and compare our hypotheses with findings from the literature. We confirm the validity of some of our hypotheses and discover new insights about the tasks that previous studies did not touch upon. We show a few examples in Table 4, and the full list of hypotheses can be found in Appendix D.
Our generated hypotheses align with useful features in existing literature.
For Deceptive reviews, we find that deceptive reviews are more likely to be emotional, use superlatives, or contain information that could not have been directly experienced. Similar findings are also found by previous studies on Deceptive reviews (Lai et al., 2020; Anderson & Simester, 2014; Ott et al., 2011; Li et al., 2014). For Tweet popularity, we discover that tweets that are concise, with specific or relevant hashtags, or with emotional tones are more likely to be retweeted more, aligning with prior studies (Tan et al., 2014; Gligorić et al., 2019). For we find that revealing something new or using vivid language and imagery can drive engagement from readers to click on headlines. Previous studies also find these rules apply to online news headlines (Banerjee & Urminsky, 2021; Sadoski et al., 2000).
@@
Deceptive reviews Deceptive reviews contain more emotional terms. Li et al. (2014)
Truthful reviews would mention weddings or special occasions. [Uncaptioned image]
Using vivid language and imagery helps. Banerjee & Urminsky (2021)
Headlines that frame the content in a personal or relatable way are clicked more. [Uncaptioned image]
Tweet popularity Tweets with emotional tones are retweeted more. Tan et al. (2014)
Mentioning influential individuals or organizations leads to more retweets. [Uncaptioned image]
@@
We also discover new insights with our generated hypotheses.
For the Deceptive reviews dataset, truthful reviews could mention the reviewer’s purpose for staying at the hotel (e.g., business trip, vacation), but deceptive ones tend not to have this information. For we find that headlines that frame the content in a personal or relatable way are clicked more. For Tweet popularity, tweets that mention influential individuals or organizations are more likely to be retweeted.
Intriguingly, one of our hypotheses contradicts a feature engineering result.
Ott et al. (2011) find that the token “future” is associated with deceptive reviews, while one of our hypotheses says that mentions of “past experiences or future travel plans” are indicative of truthfulness. This discrepancy is interesting, because the context for the token “future” is unclear. It could be in the context of future plans but could also be as a complaint about “never going to stay at the hotel in the future.” Feature engineering is limited by contextual ambiguity, whereas our generated hypotheses and their interpretation by LLMs overcome such limitations.
Our automatic evaluation of hypothesis quality also reflects negative findings.
Given mixed evidence from previous literature on the effect of “reading ease” on headline clicks, Banerjee & Urminsky (2021) finds that reading ease negatively impacts click-through rates in hrough careful feature engineering. Consistent with this result, we found that the hypotheses that claim “straightforward” and “clear” writing to be indicative of higher click-through rates have relatively lower accuracies during training.
@@ 5 Additional Related Work
Our work is connected to many recent studies on using LLMs to propose “hypotheses”, notably, Qiu et al. (2024) and Zhong et al. (2023). Qiu et al. (2024) is motivated by testing the ability of LLMs to perform human-like induction reasoning, and Zhong et al. (2023) aims to support open-ended exploration. Similar to Qiu et al. (2024), Tenenbaum et al. (2011) is motivated by human inductive reasoning and examines concept induction in synthetic settings. Ellis et al. (2020) further learns to program concepts. Yang et al. (2024a) performs LLM-based inductive reasoning with a dataset that requires existing fact-rule pairs, which is not applicable in our real-world problems. Romera-Paredes et al. (2024) generates programs that lead to mathematical discovery. Similar to Zhong et al. (2023), Pham et al. (2024) generates and refine a list of topics to achieve interpretable topic modeling for open-ended exploration. Honovich et al. (2022) explores the deduction of task description from examples. Additionally, Qi et al. (2023), Wang et al. (2024), and Baek et al. (2024) use LLMs to generate hypotheses from previous literature. Yang et al. (2024b) tries to generate hypotheses from raw web corpus, but their method is not automated or scalable as it requires human annotated hypotheses from existing literature. Our work, in contrast, focuses on hypothesis generation between the input and the label for real-world challenging tasks and uses a UCB-style reward to propose novel algorithms.
Reasoning with LLMs.
@@
Although it is not our primary goal, our results show that hypothesis-based classifiers can outperform few-shot prompting. As hypotheses may be viewed as a form of reasoning, it is related to reasoning with LLMs (Wei et al., 2022; Wang et al., 2023, i.a.). In particular, our work differs from chain-of-thought reasoning because no predefined reasoning structure is available. Moreover, an important distinction between reasoning and hypothesis generation is that the former leverages established reasoning, while the latter requires both proposition and verification of the hypotheses, to discover unknown knowledge.
LLMs for (social) sciences.
@@
Increasing attention has been brought to the use of LLMs in social science research (Ziems et al., 2024; Kim & Lee, 2023, i.a.). Our experiments demonstrate the potential of LLMs in generating hypotheses for social science research to discover unknown knowledge in the data. Furthermore, our approach can be extended to natural sciences for general scientific discovery.
6 Conclusion & Further Discussion
@@
In this work, we propose HypoGeniC, a novel data-driven and automated method that leverages LLMs to generate hypotheses with the goal of discovering unknown knowledge. With HypoGeniC, we are not only are able to generate human-interpretable hypotheses but also achieve better predictive performance against competitive baselines and even oracles. Furthermore, our method can generalize well with different models and datasets, including open models. Notably, with our generated hypotheses, we uncover new insights in real-world tasks that are widely studied in social sciences.
The key to success in HypoGeniC is not that LLMs remembers the correct hypotheses, but lies in their ability to “hallucinate” and combine potentially relevant concepts. The exploration-exploitation process then identifies the valuable hypotheses. HypoGeniC can be directly applied to complex social science tasks. We encourage future work to explore hypothesis generation that requires additional modalities and/or leverages existing literature along with past observations.
@@ 7 Limitations
We address common concerns using a Q&A format.
Q: Why only experiment with social science tasks?
A: Math and physics problems and hypotheses are hard to represent in natural language and usually require symbolic parsers (Trinh et al., 2024). We leverage LLMs to perform tasks that it is naturally adept at, which lead us to social science tasks. We find that HypoGeniC demonstrates strong results for the selected tasks, indicating new possibilities in using LLMs for scientific discovery. We leave extending our framework to natural science tasks as future work.
Q: Why is HypoGeniC effective, given that the accuracy improvement is not significant in some settings?
A: Even if there is no significant improvement in accuracy, the benefits of HypoGeniC are found in the quality of hypotheses. We find that the generated hypotheses discover new patterns that were previously unseen, as discussed in § 4.3. Additionally, it is worth noting that LLMs are imperfect at reasoning. Thus, hypothesis-based inference with LLMs may not accurately reflect the quality of the hypotheses.
Q: Since you worked on some old datasets, what if the LLMs have pre-trained knowledge about these tasks?
A: In Table 1, the zero/few-shot learning results suggest that the models cannot solve the tasks by memorizing the data. Additionally in § 4.3, we show that HypoGeniC reveal new hypotheses, based on the literature space that we can manually search. Even if the models have been pre-trained on the datasets, these hypotheses were not reported in previous literature. This suggests that even experienced researchers still struggle in finding the hypotheses that HypoGeniC generate.
Q: What hyperparameters have you tried?
A: We aim to provide a robust framework for hypothesis generation, as opposed to focusing on the optimization of results. Thus, we did not perform an extensive hyperparameter search with the generation portion of HypoGeniC. We did not adjust the value of $k$, which determines $\mathcal{H}_{\mathrm{top}}$ in Algorithm 1 to maintain efficiency. Additionally, we only considered the effect of using a hypothesis bank size of $3$ and $20$ to only test using an extremely small hypothesis bank size and a large one. The ideal hypothesis bank size may require further investigation. Finally, we only tested the size of our wrong example bank $w_{max}$ as $10$ to strike a balance between context window sizes and generation of good quality hypotheses. We believe that a more thorough hyperparameter search could improve the performance of our methodology.
Q: How costly is your approach?
A: HypoGeniC has high latency, specifically when using inference methods that require multiple prompts. For example, the filter and weighted vote inference policy requires iterating through the top hypotheses to determine relevance and then performing inference if it is relevant. For single-step adaptive inference and best accuracy hypothesis, however, HypoGeniC is efficient. Given that we request reasoning for all inference prompts, the procedure can be time-consuming and require financial costs (e.g., GPT-3.5-turbo takes \$2.05 on average over 76 experiments with an average of 1.5 hours per experiment). This concern is alleviated when using open models. However, all these processes are still relatively cheap compared to human efforts.
Q: What are some potential risks of hypothesis generation?
A: One potential risk of hypothesis generation is that there is little guard regarding steorotypes and biases being confirmed if given data that may seem to enforce them. As a result, it can be potentially harmful to use HypoGeniC in a real-world setting without proper oversight. Additionally, if the data reveals personal information regarding people, there is no guarantee that the hypotheses generated will not reveal this information. We highly recommend human-AI collaboration in using HypoGeniC to ensure that the generated hypotheses are ethical and unbiased.
@@ A.4 Retweeted More
Appendix B Implementation and Setup Details
@@ B.3 Hyperparameters
B.4 Licensing Details
The Deceptive reviews and Tweet popularity datasets have not been released with any licenses, but are free to use for research purposes based upon the authors. The ataset is released under the Creative Commons Attribution 4.0 International License. The Shoe sales dataset will be released under the same licensing as this work, CC BY 4.0 License, should it be accepted.
In regards to models, we find that GPT-3.5-turbo and Claude-2.1 are all proprietary models and are not released under any open-source licenses. On the other hand, Mixtral is released under the Apache License 2.0. RoBERTa is not released under specific licensing but is free to use for research purposes. However, Llama-2-7B is released under their own licensing found at https://ai.meta.com/llama/license/.
Per our extensive search, we find that we are in compliance with the licensing agreements of all the datasets and models used in this work.
@@ C.1 HypoGeniC Performance across inference strategies
We also present our results with confidence intervals. We specifically see that compared to the Oracle Methods, HypoGeniC shows performance statistically significant benefits when comparing to the 200 training examples for nd Tweet popularity. However, this is not the case for Deceptive reviews, because there are word level features that make the task easier for unsupervised methods. We note that HypoGeniC has statistically significant performance increases for Deceptive reviews with Claude-2.1 and Mixtral and for Tweet popularity with Claude-2.1 and Mixtral.
Shoe Deceptive Headline Tweet
Models Methods Sales Reviews Popularity Popularity
RoBERTa (Oracle) Train 200 100.0 $\pm$ 0.0 84.0 $\pm$ 4.2 49.0 $\pm$ 5.7 50.7 $\pm$ 5.7
Train 1000 100.0 $\pm$ 0.0 91.0 $\pm$ 3.2 60.0 $\pm$ 5.5 62.0 $\pm$ 5.5
Llama-2-7B (Oracle) Train 200 100.0 $\pm$ 0.0 88.7 $\pm$ 3.6 49.7 $\pm$ 5.7 50.3 $\pm$ 5.7
Train 1000 100.0 $\pm$ 0.0 92.3 $\pm$ 3.0 60.0 $\pm$ 5.5 51.3 $\pm$ 5.7
Claude-2.1 Few shot 75.0 $\pm$ 4.9 51.0 $\pm$ 5.7 60.0 $\pm$ 5.5 0.3 $\pm$ 0.6
HypoGeniC 100.0 $\pm$ 0.0 75.3 $\pm$ 4.9 61.3 $\pm$ 5.5 62.0 $\pm$ 5.5
Mixtral Few shot 79.0 $\pm$ 4.6 56.3 $\pm$ 5.6 55.3 $\pm$ 5.6 48.7 $\pm$ 5.7
HypoGeniC 98.0 $\pm$ 1.6 68.0 $\pm$ 5.3 60.3 $\pm$ 5.5 62.7 $\pm$ 5.5
GPT-3.5-turbo Few shot 49.0 $\pm$ 5.7 55.0 $\pm$ 5.6 60.0 $\pm$ 5.5 62.0$\pm$ 5.5
HypoGeniC 100.0 $\pm$ 0.0 60.7 $\pm$ 5.5 63.7 $\pm$ 5.4 61.0 $\pm$ 5.5
Table 5: Table with 95% confidence interval for Few shot results and HypoGeniC for our best results.
C.2 HypoGeniC Performance across training examples
Figure 3: Claude-2.1 results for baselines, HypoGeniC (no update), and HypoGeniC (best) with hypothesis bank size 3 and 20 across multiple training samples
Figure 3 presents the results for the performance of HypoGeniC with Claude-2.1 as the training examples change. We observe that for all of our datasets, HypoGeniC outperforms zero-shot and few-shot learning generally for all training examples in Shoe sales and Tweet popularity. In we find that the model needs to use 200 examples to outperform them. We highlight that HypoGeniC outperforms the No Updates method for all training examples across the four datasets when using a hypothesis bank size of 20. When using a hypothesis bank size of 3, we find that in Tweet popularity, HypoGeniC is able to outperform the No Updates method, but is unable to as the training examples increase. In Shoe sales we observe that it is largely worse because we set $k$ (as discussed in § 2.1) to be 1, which causes difficulty in finding the best hypothesis. It is unclear what the optimal number of training examples is across the datasets, as using more examples does not necessarily increase accuracy.
Figure 4: GPT-3.5-turbo results for baselines, HypoGeniC (no update), and HypoGeniC (best) with hypothesis bank size 3 and 20 across multiple training samples
Figure 4 displays the accuracy for HypoGeniC with GPT-3.5-turbo for the different training examples. We observe that unlike HypoGeniC performance with Claude-2.1, our results are mixed for when our method outperforms the few shot inference. Specifically, in Tweet popularity, the few shot inference surpasses our results, indicating that in this set hypotheses provide less benefits than using examples. As HypoGeniC exceeds the accuracy of zero shot’s, the proposed method still provides benefits to the base model. Similar to the results on Claude-2.1, we outperform RoBERTa and Llama-2-7B on all datasets aside on Deceptive reviews for all training examples. HypoGeniC surpasses the performance of the No Update strategy generally for all training examples. We note that due to the limited context window of GPT-3.5-turbo, the No Update strategy fails as it is unable to accept training examples. HypoGeniC effectively bypasses this issue by iteratively going through test examples, as opposed to feeding them into the model all at once.
Figure 5: Mixtral results for baselines, HypoGeniC (no update), and HypoGeniC (best) with hypothesis bank size 3 and 20 across multiple training samples
In, Figure 5, the performance of HypoGeniC for varying training examples with Mixtral is shown. HypoGeniC outperforms the zero shot and few shot strategies for all datasets, aside from Shoe sales, where the proposed method requires 200 examples to outperform few shot learning. Similary, we note that HypoGeniC surpasses the performance of RoBERTa and Llama-2-7B for Tweet popularity, and generally for Shoe sales. As mentioned in § C.1, despite Mixtral finding the best hypothesis, it occasionally refuses to choose the correct label to encourage “variety”, which causes RoBERTa and Llama-2-7B to outpeform HypoGeniC. In comparison to the No Update results, we find that in Deceptive reviews and HypoGeniC matches or exceeds this method. For Shoe sales, we find that with hypothesis bank 3, HypoGeniC must use 200 examples, to finally converge to the correct hypothesis. On the other hand, for Tweet popularity, No Update surpasses the HypoGeniC with hypothesis bank size 3 after using 200 training examples. This may occur as using 3 hypotheses is too limited to adeptly describe the dataset, causing accuracy to suffer.
C.3 Full OOD results
Models Methods IND Deceptive Reviews OOD Deceptive Reviews
RoBERTa (Oracle) Train 200 84.0 73.0 ($\downarrow$11.011.011.011.0)
Train 1000 91.0 79.7 ($\downarrow$11.311.311.311.3)
Llama-2-7B (Oracle) Train 200 88.7 78.7 ($\downarrow$10.010.010.010.0)
Train 1000 92.3 88.7 ($\downarrow$3.63.63.63.6)
Claude-2.1 Zero shot 31.0 27.7 ($\downarrow$3.33.33.33.3)
Few shot 51.0 41.7 ($\downarrow$9.39.39.39.3)
HypoGeniC (Best-accuracy hypothesis) 67.3 71.7 ($\uparrow$4.44.44.44.4)
HypoGeniC (Filter and weighted vote) 68.0 74.7 ($\uparrow$6.76.76.76.7)
HypoGeniC (One-step adaptive) 70.0 68.3 ($\downarrow$1.71.71.71.7)
HypoGeniC (Two-step adaptive) 67.7 70.7 ($\uparrow$3.03.03.03.0)
Mixtral Zero shot 55.0 49.7 ($\downarrow$5.35.35.35.3)
Few shot 56.3 49.0 ($\downarrow$7.37.37.37.3)
HypoGeniC (Best-accuracy hypothesis) 61.3 64.7 ($\uparrow$3.43.43.43.4)
HypoGeniC (Filter and weighted vote) 62.0 61.0 ($\downarrow$1.01.01.01.0)
HypoGeniC (One-step adaptive) 63.0 54.7 ($\downarrow$8.38.38.38.3)
HypoGeniC (Two-step adaptive) 61.3 64.7 ($\uparrow$3.43.43.43.4)
GPT-3.5-turbo Zero shot 50.0 49.0 ($\downarrow$1.01.01.01.0)
Few shot 55.0 52.0 ($\downarrow$3.03.03.03.0)
HypoGeniC (Best-accuracy hypothesis) 57.3 60.7 ($\uparrow$3.43.43.43.4)
HypoGeniC (Filter and weighted vote) 55.3 55.7 ($\uparrow$0.40.40.40.4)
HypoGeniC (One-step adaptive) 55.7 51.7 ($\downarrow$4.04.04.04.0)
HypoGeniC (Two-step adaptive) 54.7 59.0 ($\uparrow$4.34.34.34.3)
Table 6: Performance of baselines and compared to our methods on the out-of-distribution deceptive reviews and Deceptive reviews.
Table 6 shows results for the OOD deceptive reviews dataset for all inference strategies for each model.
We find that HypoGeniC outperforms both zero shot and few shot learning across all models and inference policies. The best-accuracy hypothesis and two-step adaptive inference methods are the most robust, showing an average increase of 3.7% and 3.6% respectively. We claim that although the filter and weighted vote strategy at first glance may seem to have mixed performance, the method is still robust. The drop in accuracy for Mixtral with filter and weighted is minimal (1%), and both GPT-3.5-turbo and Claude-2.1 exhibit increases in accuracy. Hence, the inference policy is consistent across Deceptive reviews and the OOD deceptive review datset. Interestingly, the single-step adaptive inference method exhibits drops in performance despite being the best performing inference model in Deceptive reviews. In single-step adaptive inference, the LLM sees both the hypotheses with the sets of examples along with the final question of determining whether the review is deceptive. Even though the LLM is prompted to only use one chosen hypotheses, these training examples from Deceptive reviews negatively impact the model because they are part of the context and are thus inherently used by LLMs. On the other hand, for two-step adaptive inference, since there is a dedicated prompt for hypothesis selection, the application of the hypothesis is unaffected from the Deceptive reviews training samples.
Appendix D Qualitative Analysis on Generated Hypotheses
We include findings from the generated hypotheses on Deceptive reviews, and Tweet popularity datasets in Table 7. The table shows that the a good number of the hypotheses are supported by existing findings, while others are novel. This suggests that the generated hypotheses are grounded in existing literature and can be used to guide future research.
Dataset Finding Supported/Novel
Deceptive reviews Deceptive reviews contain more emotional terms. Li et al. (2014)
Deceptive reviews are more likely to use superlatives. Ott et al. (2011)
Deceptive reviews contain hearsay or information that could not have been directly experienced. Ott et al. (2011)
Deceptive reviews tend to be more exaggerated. Anderson & Simester (2014)
Truthful reviews tend to use more balanced and objective tone. Anderson & Simester (2014)
Truthful reviews could mention the reviewer’s purpose for staying at the hotel (e.g., business trip, vacation). Novel
Truthful reviews would mention weddings or special occasions. Novel
Truthful reviews may contain information about reviewer’s expectations and previous hotel experiences. Novel
Truthful reviews would acknowledge the reviewer’s personal biases or preferences. Novel
Deceptive ones may present the reviewer’s opinion as objective facts. Novel
Truthful reviews may contain reviewers’ past experiences or future travel plans. Novel
Concreteness helps. Sadoski et al. (2000)
Revealing something new helps. Banerjee & Urminsky (2021)
Using vivid language and imagery helps. Banerjee & Urminsky (2021)
Headlines with high intensity of emotions would be clicked more. Banerjee & Urminsky (2021)
Action-oriented headlines are clicked more. Banerjee & Urminsky (2021)
Humorous headlines are clicked more. Novel
Controversial headlines are clicked more. Novel
Headlines that frame the content in a personal or relatable way are clicked more. Novel
Tweet popularity Short and concise tweets are retweeted more. Gligorić et al. (2019)
Tweets with emotional tones are retweeted more. Tan et al. (2014)
Including specific details (e.g., dates, locations) are associated with more retweets. Novel
Including statistics and data are associated with more retweets. Novel
Mentioning influential individuals or organizations leads to more retweets. Novel
Including links to additional content (e.g., articles, videos) leads to more retweets. Novel
Tweets with a call to action or urgency are found to be retweeted more. Novel
Table 7: Summary of generated hypotheses (on the real-world datasets) and whether they support existing findings or are novel.